# ENFORCE: Exact Nonlinear Constrained Learning with Adaptive-depth Neural Projection

### v2

@@ 4.1 AdaNP: Adaptive-depth neural projection
While the KKT conditions for an NLP (Eq. 1) can be more generally solved using Newton’s methods, our method circumvents the computational overhead associated with calculating the Hessian matrix of the constraints (Appendix A.1).
@@ 4.2 Architecture
The architecture of ENFORCE (Fig. 1) is composed of (1) a backbone neural network, that can be of any kind and complexity, and (2) an AdaNP module. The depth of AdaNP depends on the backbone performance and specified tolerance. Indeed, AdaNP can be purposely tuned to increase training efficiency (Section 4.4). A single NP layer is composed of two steps: (1) automatic differentiation and (2) local neural projection.
@@
To compute the Jacobian of the constraint system, we leverage automatic differentiation available in most deep learning libraries [48, 49, 50]. Computing the local Jacobian $J_{y}c|_{x,\hat{y}}$ is computationally inexpensive, as it requires propagating derivatives only through the constraints and does not involve the neural network backbone. Furthermore, its computation can be efficiently parallelized on GPU.
@@
The neural projection defined by the operator $\mathcal{P}$ in Def. 1 depends on individual input-prediction instances. Thus, the projection is locally defined in the neighborhood of ($x_{i},\hat{y}_{i}$). We parallelize the computation of local neural projections by building a rank-3 tensor $\mathbf{B}^{*}$ and a rank-2 tensor $\mathbf{v}^{*}$ (Appendix A.2). Thus, we reduce the cost of an NP layer from $\mathcal{O}(BS\times N_{C}^{3})$ to $\mathcal{O}(N_{C}^{3})$, allowing effective training with stochastic gradient descent techniques with virtually no limitation on the batch dimension ($BS$). For modern hardware, handling output dimensionalities of $N_{C}<10^{3}$ leads to minimal computational cost.
4.3 Implications of orthogonal projections
The set of constraints $c(x,y)=0$ describes an infinite-wide feasible region (i.e., a hypersurface) where the constraints are defined. Hence, one could ask whether the projections are unique and thus suitable for guiding the learning process. To this end, theoretical insights can be derived by analyzing the case of orthogonal projections, i.e., $W=I$ (Eq. 1).
(a) Affine constraint
(b) Nonlinear constraint
Figure 2: Orthogonal projection on the feasible region when $y\in\mathbb{R}^{2}$. Orthogonal projections on affine constraints are unique and always lead to enhanced predictions. When the constraint is nonlinear, multiple orthogonal projections are possible, not necessarily improving the prediction.
Affine constraints
Considering constraints that are affine in $y$, the feasible space is a hyperplane defined in $\mathbb{R}^{N_{O}}$. Under those circumstances, it can be proven that there exists a unique orthogonal projection defining $\tilde{y}$ from a point $\hat{y}\in\mathbb{R}^{N_{O}}$ such that $c(\tilde{y})=0$. More interestingly, if $y^{*}$ is the ground truth, then $\forall\hat{y},y^{*},\,||y^{*}-\tilde{y}||<||y^{*}-\hat{y}||$, that is, the projected prediction is always a better prediction than the original one (Fig. 2(a)). In such a condition, constrained learning effectively guides the learning process toward the ground truth. In particular, the loss is expected to decrease compared to the unconstrained case from the very start of the training process.
Nonlinear constraints
The same considerations do not hold for nonlinear hypersurfaces. Firstly, there could be multiple points on the hypersurface where the normal vectors point toward the prediction $\hat{y}$. Secondly, the closest orthogonal projection can potentially represent a worse prediction, since the distance $||y^{*}-\tilde{y}||$ is not ensured to be shorter than $||y^{*}-\hat{y}||$ (see Fig. 2(b)). Then, the projection step could potentially conflict with the training step, preventing the training process itself from converging. Therefore, we can not conclude that orthogonal projections onto a nonlinear feasible region always guide the learning process efficiently.
@@ 4.4 Training ENFORCE
We train ENFORCE using standard unconstrained gradient descent methods (i.e., Adam). We develop and use a constrained learning methodology using AdaNP to guide the neural network training to convergence, supported by the theoretical implications of orthogonal projections described above (Section 4.3).
Linearized constraints and training strategy
ENFORCE does not project the predictions directly on the nonlinear feasible region but on a locally linearized counterpart. Thus, we can assume the existence of a unique closest orthogonal projection. However, the quality of the latter, i.e., whether the projection really improves the prediction, is largely influenced by the constraint geometry curvature. Moreover, the initialization of the neural network plays a crucial role in constrained learning using ENFORCE. Overall, it is difficult to prove a priori whether constrained learning can be beneficial during the initial phase of training.
To overcome this limitation, we develop a strategy to ensure effective constrained learning. Instead of building a static ENFORCE model, we activate AdaNP only when its application enhances prediction accuracy (i.e., the loss decreases). This strategy can be interpreted as an unconstrained pre-training; however, it is automatically incorporated in ENFORCE. We have demonstrated that this approach accelerates the training process and leads to better optima compared to fully unconstrained training.
@@
We further control the constrained learning process by minimizing the projection displacement, i.e., the distance between the preliminary NN prediction $\hat{y}$ and its projection $\tilde{y}=\text{AdaNP}(\hat{y})$. This aims to (1) ensure minor linearization error ($\epsilon_{L}\sim\Delta y$) and (2) prevent the neural network from learning alternative functions whose projections fall within the neighborhood of the desired functions. The latter is suggested to reduce reliance on AdaNP for accurate predictions, thereby lowering the computational cost during inference (i.e., by decreasing the depth of AdaNP).
We control the projection displacement by applying soft constraints, adding an extra penalty term to the loss function:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}||y_{i}-\tilde{y}_{i}||^{2}+\frac{\lambda}{N}\sum_{i=1}^{N}||\hat{y}_{i}-\tilde{y}_{i}||^{2}$ (4)
Here, the first term represents the conventional mean squared error between the projected prediction $\tilde{\mathbf{y}}_{i}$ and the ground truth $\mathbf{y}_{i}$, while the second term quantifies the projection displacement, targeted at controlling and minimizing its magnitude, by acting as a penalty. The two terms are balanced by a scalar weight $\lambda$.
@@ 5 Experiments and discussion
We demonstrate our proposed method on an illustrative two-dimensional example. We aim to fit the oscillating function $y:\mathbb{R}\rightarrow\mathbb{R}^{2}$:
@@
where $x\in\mathbb{R}$ is the unidimensional independent variable (input) and the scalar $f$ is the frequency. Notably, the system is implicitly linked by a nonlinear constraint $c(x,y_{1},y_{2})=(0.5y_{1})^{2}+x^{2}+y_{2}$, involving both input and output variables. We train an ENFORCE model consisting of a 64-neuron 1-hidden layer fully connected ReLU neural network as a backbone and an AdaNP module to force the predictions to satisfy the constraint. We experiment with different constraint tolerances ($\epsilon_{T}$) and loss weighting factors ($\lambda$) during training to heuristically study their influence. At inference, the depth of AdaNP is adapted to guarantee a specified tolerance $\epsilon$, e.g., absolute average or maximum residual across a batch of prediction. In this case study, we set the average absolute residual to be below $\epsilon_{I}=10^{-6}$ at inference time. To verify the interpolation capabilities, we sample 100 training data points from a uniform distribution in $x=[-2,2]$ and 1000 test points in the same domain. Every run is repeated 5 times using different initialization seeds.
Accuracy and constraint guarantee
One of the major concerns in constrained learning is related to keeping the expressivity and approximation capabilities of NNs [51, 52] while enforcing constraints to be satisfied. Here, we evaluate the regression capabilities and robustness to spectral biases [53] of ENFORCE on oscillating functions ($f=5$) while assessing the average and maximum constraint deviation (i.e., residual). We compare the method with a traditional (unconstrained) multilayer perceptron (MLP) and a soft-constrained neural network sharing the same architecture. In the soft-constrained neural network, the constraint residual is added as an additional loss term balanced with a scalar weight ($\lambda_{y}$). We report here the outcomes of the best-performing soft-constrained network, with $\lambda_{y}=1$. We train all the NNs for 50,000 epochs, using Adam [54] optimizer with a learning rate of $0.001$.
The main results are summarized in Table 1. ENFORCE outperforms the soft-constrained NN and the MLP by effectively minimizing the nonlinear constraint residual, guaranteeing arbitrary satisfaction with minor computational costs (Fig. 3(c)). To quantify the relative impact, we normalized the residuals by scaling them with respect to the absolute maximum value of the variables within the considered range ($\max(|x|,|y_{1}|,|y_{2}|)$). It is worth noting how overall good performance (e.g., in NN) does not guarantee compliance with underlying model constraints, with deviations above 17%. By contrast, ENFORCE demonstrates a worst-case residual below 0.0001% (Table 1). The inference time for a batch of 1000 samples is 6 ms longer when using ENFORCE compared to an MLP. This amount should be regarded as additive, not multiplicative (e.g., 4x relative to the MLP). Indeed, the computational complexity is entirely attributed to the AdaNP module, meaning that the backbone architecture has no impact. Therefore, when applied to larger models (e.g., transformers), the relative computational impact becomes negligible. Indeed, the complexity is primarily influenced by the number of constraints and the batch size, with the latter being particularly significant when inference is performed on CPU hardware (see Section 4.1 and Appendix A.2).
Besides constraint guarantee, ENFORCE also enhances the accuracy of the MLP (the $R^{2}$ increases from 0.994 to 0.999). At first sight, this may appear to be only a marginal improvement. However, advancing deep learning methods to achieve near-perfect performance remains an open challenge in essential fields such as healthcare and self-driving cars, where even slight inaccuracies at the decimal level can lead to catastrophic consequences [3, 4].
Interestingly, ENFORCE outperforms the MLP in data-scarce regions of the domain, which in this dataset correspond to the domain extremities (as shown in Fig. 3(b)). More generally, ENFORCE also performs better under data-scarcity conditions when the models are trained on uniformly sampled fractions of the dataset (Fig. 3(d)). This observation suggests that constrained learning may enhance data efficiency.
@@
Table 1: Regression accuracy and constraint guarantee of ENFORCE when compared with a multilayer perceptron (MLP) and a soft-constrained NN (Soft). Results for $\lambda=0.5$, $\epsilon_{T}=10^{-4}$, and $\epsilon_{I}=10^{-6}$ are reported. The inference on 1000 samples is run on a CPU (11th Gen Intel(R) Core(TM) i7-1185G7 @ 3.00GHz, 4 Core(s)) to prove efficiency also in the absence of GPU hardware. The inference time is also reported for a batch of 1000 predictions.
MAPE [%] $R^{2}$ Avg Residual [%] Max Residual [%] Inference Time [s]

### v3

@@ 4.1 AdaNP: Adaptive-depth neural projection
While the KKT conditions for a nonlinear program (Eq. 1) can be more generally solved using Newton’s methods, our method circumvents the computational overhead associated with calculating the Hessian matrix of the constraints (cf. Appendix B.2) at the cost of reduced convergence rate (i.e., full Newton’s method converges quadratically).
4.2 Conditioning analysis
The set of constraints $c(x,y)=0$ describes an infinite-wide feasible region (i.e., a hypersurface) where the constraints are defined. Hence, one could ask whether the projections are unique as well as whether the projection mapping is stable and well-conditioned, and thus suitable for guiding the learning process. We provide theoretical conditions for the uniqueness and no-worse property of the projected prediction, along with evidence of regularity and robustness, and show that other state-of-the-art constrained learning approaches do not guarantee these properties.
Proposition 2 (Uniqueness of the projection).
We consider the submanifold defined by the linearly independent constraints in the output space =$\mathcal{M}=\{y\in\mathbb{R}^{N_{O}}:c(y)=0\}$, where $c$ are real-valued, continuously differentiable functions with Lipschitz-continuous gradients (hence, the manifold has positive reach [35]). If $\hat{y}\in\operatorname{reach}(\mathcal{M})$, then there exists a unique orthogonal projection of $\hat{y}$ onto $\mathcal{M}$ (cf. Appendix B.4 and [35]).
The reach depends on the geometry of the constraint set (e.g., the reach is infinite for linear constraints and approaches zero near sharp corners or singularities). Therefore, for the projection to be well-defined, $\hat{y}$ must lie within a sufficiently small tubular neighborhood of $\mathcal{M}$.
Proposition 3 (No-worse property).
When a unique projection $\tilde{y}$ exists and the manifold $\mathcal{M}$ has positive reach, then the projection mapping satisfies the no-worse property: for any prediction $\hat{y}$ and ground truth $y^{*}$, $||y^{*}-\tilde{y}||\leq||y^{*}-\hat{y}||$, that is, the projected prediction is always a no-worse prediction than the original one (cf. Appendix B.5).
Regularity and robustness of appended layers (e.g., projection operation or null-space completion) can influence the stability of the neural network training. We prove (cf. Appendix B.6) the projection operation to be a non-expansive mapping in the neighborhood of the constraint manifold (i.e., its Lipschitz constant $L\leq 1$). This ensures adversarial robustness of the neural projection layer during the forward pass and stable gradient flow dynamics.
Theorem 1 (Non-expansiveness of the projection operator).
Given the nonlinear program:
$\tilde{y}=\operatorname*{arg\,min}_{y}\frac{1}{2}||y-\hat{y}||^{2}\quad\text{s.t.}\quad c(y)=0,$ (4)
with $\hat{y}\in\mathbb{R}^{N_{O}}$, $\tilde{y}\in\mathbb{R}^{N_{O}}$, and $c(y):\mathbb{R}^{N_{O}}\rightarrow\mathbb{R}^{N_{C}}$ smooth, continuous constraints with a full-rank Jacobian $J_{y}c$, the projection operator $\mathcal{P}:\tilde{y}=\mathcal{P}(\hat{y})$ solving the nonlinear program (Eq. 4) is a non-expansive mapping in the neighborhood of the constraints manifold:
$\forall\hat{y}_{i},\hat{y}_{j}\in\mathbb{R}^{N_{O}},$ $\operatorname{dist}(\hat{y}_{i},\mathcal{M})<\operatorname{reach}(\mathcal{M})\land\operatorname{dist}(\hat{y}_{j},\mathcal{M})<\operatorname{reach}(\mathcal{M})$ (5)
$\Rightarrow\ \left\|\mathcal{P}(\hat{y}_{i})-\mathcal{P}(\hat{y}_{j})\right\|\leq\left\|\hat{y}_{i}-\hat{y}_{j}\right\|$
Lemma 1 (Gradient flow dynamics).
The non-expansive property of the projection mapping ensures the stability of the gradient flow dynamics. During backpropagation, the pre-projection gradient is multiplied by a factor $J_{\hat{y}}\mathcal{P}^{T}$, which is the bounded Lipschitz constant. If $\nabla_{\tilde{y}}\ell$ is the post-projection gradient, then we can bound the pre-projection gradient:
$\nabla_{\hat{y}}\ell=J_{\hat{y}}\mathcal{P}^{T}\nabla_{\tilde{y}}\ell\>\Rightarrow\>||\nabla_{\hat{y}}\ell||\leq||J_{\hat{y}}\mathcal{P}^{T}||\,||\nabla_{\tilde{y}}\ell||$ (6)
Thus, no exploding gradients can originate from the projection step, as the scale of pre-projection and post-projection gradients is comparable. Conversely, alternative constrained learning paradigms do not guarantee bounds on the operated mapping. For example, predict-and-complete approaches using Newton-based root-finding methods (e.g., DC3 [7]) can exhibit non-Lipschitz behavior when the constraint Jacobian is ill-conditioned (cf. Appendix B.6), potentially leading to unstable training and slower convergence.
@@ 4.3 Architecture
The architecture of ENFORCE (Fig. 1) is composed of (1) a neural network (without loss of generality) as backbone, which can be of any kind and complexity, and (2) an AdaNP module. The depth of AdaNP depends on the backbone performance and specified tolerance. Indeed, the tolerance of AdaNP can be tuned to increase training efficiency (cf. Appendix D.1). A single NP layer is composed of two steps: (1) automatic differentiation and (2) local neural projection.
@@
To compute the Jacobian of the constraint system, when not available analytically, we leverage automatic differentiation available in most deep learning libraries [36, 37, 38]. Computing the local Jacobian $J_{y}c|_{x,\hat{y}}$ is computationally inexpensive, as it requires propagating derivatives only through the constraints and does not involve the neural network backbone. Furthermore, its computation can be efficiently parallelized on GPU.
@@
The neural projection defined by the operator $\mathcal{P}$ in Def. 1 depends on individual input-prediction instances. Thus, the projection is locally defined in the neighborhood of ($x_{i},\hat{y}_{i}$). We parallelize the computation of local neural projections by building a rank-3 tensor $\mathbf{B}$ and a rank-2 tensor $\mathbf{v}$ (Appendix D.3). Thus, we reduce the apparent complexity of an NP layer from $\mathcal{O}(BS\times N_{C}^{3})$ to $\mathcal{O}(N_{C}^{3})$, allowing effective training with stochastic gradient descent techniques. On modern hardware, handling up to $N_{C}<10^{3}$ constraints results in a computational cost that remains practical, particularly when using Cholesky decomposition for matrix inversion [39]. Moreover, the complexity of this method is equivalent to other state-of-the-art methods such as DC3 [7]. Additional insights into memory requirements are provided in Appendix D.4.
@@ 4.4 Training ENFORCE
We train ENFORCE using standard unconstrained gradient descent methods (i.e., Adam [32]). We develop and use a constrained learning methodology using AdaNP to guide the neural network training to convergence, supported by the theoretical implications of orthogonal projections described in Section 4.2.
@@
The loss function used throughout this study takes the following general form:
$\ell=\ell_{T}+\ell_{D}+\ell_{C}=\ell_{T}+\frac{\lambda_{D}}{N}\sum_{i=1}^{N}||\hat{y}_{i}-\tilde{y}_{i}||^{2}+\frac{\lambda_{C}}{N}\sum_{i=1}^{N}||c(x_{i},\tilde{y}_{i})||,$ (7)
where the first term, $\ell_{T}$, is a task-specific loss function selected based on the target model. The second and third terms are regularization penalties that respectively minimize the projection displacement, $||\hat{y}_{i}-\tilde{y}_{i}||$, and the constraint residual $||c(x_{i},\tilde{y}_{i})||$. The relative contributions of these terms are controlled by the scalar weights $\lambda_{D}$ and $\lambda_{C}$. Minimizing the projection displacement aims to (1) ensure minor linearization error ($\epsilon_{L}\sim\Delta y$) and (2) prevent the neural network from learning alternative functions whose projections onto the constraints fall within the neighborhood of the desired functions. Also, this additional loss term is suggested to reduce reliance on AdaNP, thereby lowering the computational cost during inference (i.e., by decreasing the depth of AdaNP).
Adaptive training strategy
We propose a strategy to facilitate constrained learning during the early stages of training, guided by the theoretical insights presented in Section 4.2. In the initial training phases, the preliminary prediction $\hat{y}$ may be inaccurate and lie far from the constraint manifold $\mathcal{M}$. Under such conditions, projecting onto a locally linearized approximation of the constraints can introduce substantial errors in the prediction. To mitigate this issue in practice, inspired by trust-region methods [33], we activate AdaNP only when the projection operation leads to an improvement in the prediction accuracy (e.g., quantified by a decrease in some loss measure $m_{\ell}$). This often leads to an unconstrained pre-training phase, followed by the activation of the AdaNP module. In other words, this serves as a heuristic to ensure that the prediction $\hat{y}$ lies sufficiently close to the constraint manifold $\mathcal{M}$. Details about the underlying algorithm and loss measures used in this study are reported in Appendix D.2.
@@ 5 Experiments and discussion
We evaluate the proposed method on two tasks: (i) fitting a vector-valued function with implicitly and nonlinearly related components, and (ii) learning solutions to nonlinear parametric optimization problems. While the function fitting case study is kept low-dimensional for illustrative purposes, we benchmark ENFORCE on high-dimensional optimization problems following the state-of-the-art experimental routine in DC3 [7]. All experiments were conducted using an NVIDIA A100 Tensor Core GPU 80 GB, while the nonlinear programming solver (Section 5.2) runs on a CPU (11th Gen Intel(R) Core(TM) i7, 4 Core(s), 8 Logical Processor(s)).
5.1 Illustrative function fitting
We aim to fit the illustrative oscillating function $y:\mathbb{R}\rightarrow\mathbb{R}^{2}$:
@@ 5.1 Illustrative function fitting
where $x\in\mathbb{R}$ is the unidimensional independent variable (input) and the scalar $f=5$ is the frequency. Notably, the system is implicitly linked by a nonlinear constraint $c(x,y_{1},y_{2})=(0.5y_{1})^{2}+x^{2}+y_{2}$, involving both input and output variables. We train an ENFORCE model consisting of a 64-neuron 1-hidden-layer fully connected ReLU neural network as a backbone and an AdaNP module to force the predictions to satisfy the constraint. The supervised task loss is the mean squared error ($\ell_{T}=\frac{1}{N}\sum_{i=1}^{N}||y_{i}-\tilde{y}_{i}||^{2}$), while $\lambda_{C}$ is set to zero (i.e., the constraint is addressed exclusively by AdaNP and no soft constraint term is used). At inference time, the depth of AdaNP is adaptively adjusted to enforce the constraint within a specified tolerance $\epsilon$, such as the absolute average or maximum residual across a batch of predictions. In this case study, we set the maximum absolute residual to be below $\epsilon_{I}=10^{-6}$ at inference time. To verify the regression capabilities, we sample 100 training data points from a uniform distribution in $x=[-2,2]$ and 100,000 test points in the same domain. We compare the method with a multilayer perceptron (MLP) and a soft-constrained neural network sharing the same architecture. We report here the outcomes of the best-performing soft-constrained network, with a constraint loss weight $\lambda_{C}=1$. We train all the models for 50,000 epochs, using Adam optimizer [32] and a learning rate of $10^{-3}$. Every run is repeated 5 times using different initialization seeds.
@@
Table 1: Regression accuracy and constraint guarantee of ENFORCE on 100,000 test samples when compared with a multilayer perceptron (MLP) and a soft-constrained neural network (Soft). Results for $\lambda_{D}=0.5$, $\epsilon_{T}=10^{-4}$, and $\epsilon_{I}=10^{-6}$ are reported. We report the inference time for a batch of 1,000 samples. (Note that $\text{MAPE}=\frac{100\%}{N}\sum_{i=1}^{N}\left|\frac{y^{*}_{i}-\tilde{y}_{i}}{y^{*}_{i}}\right|$).
Method MAPE [%] $R^{2}$ Mean eq. [%] Max eq. [%] Inf. time [s]